# Interpreting Conversational Dense Retrieval by Rewriting-Enhanced Inversion of Session Embedding

### v1

@@ Abstract
Conversational dense retrieval has shown to be effective in conversational search. However, a major limitation of conversational dense retrieval is their lack of interpretability, hindering intuitive understanding of model behaviors for targeted improvements. This paper presents ConvInv , a simple yet effective approach to shed light on interpretable conversational dense retrieval models. ConvInv transforms opaque conversational session embeddings into explicitly interpretable text while faithfully maintaining their original retrieval performance as much as possible. Such transformation is achieved by training a recently proposed Vec2Text model Morris et al. (2023) based on the ad-hoc query encoder, leveraging the fact that the session and query embeddings share the same space in existing conversational dense retrieval. To further enhance interpretability, we propose to incorporate external interpretable query rewrites into the transformation process. Extensive evaluations on three conversational search benchmarks demonstrate that ConvInv can yield more interpretable text and faithfully preserve original retrieval performance than baselines. Our work connects opaque session embeddings with transparent query rewriting, paving the way toward trustworthy conversational search. Our code is available at this anonymous repository.
@@
[…]
@@ 1 Introduction
With the rapid development of language modeling, conversational search has emerged as a novel search paradigm and is garnering more and more attention. Different from the traditional ad-hoc search paradigm characterized by keyword-based queries and “ten-blue” links (Yu et al., 2020), conversational search empowers users to interact with the search engine through multi-turn natural language conversations to seek information, which brings a more intuitive and efficient search experience (Mao et al., 2022b; Gao et al., 2022).
In conversational search, the system input is a multi-turn natural language conversation, which may have many linguistic problems such as omissions, co-references, and ambiguities (Radlinski and Craswell, 2017), posing great challenges for accurately grasping the user’s real information needs. Recently, conversational dense retrieval (CDR) (Yu et al., 2021; Lin et al., 2021; Kim and Kim, 2022; Mao et al., 2022a; Qian and Dou, 2022; Mo et al., 2023b), which directly encodes the whole conversational search session and the passages into a unified embedding space to perform matching, has shown to be a promising method to solve this complex search task. Compared to another type of method: conversational query rewriting (CQR) (Lin et al., 2020; Vakulenko et al., 2021a; Wu et al., 2022; Mo et al., 2023a), which is a two-step method that first reformulates the search session into a decontextualized query rewrite and subsequently inputs this rewrite into existing ad-hoc search models for search, the end-to-end CDR models can be directly optimized towards better search effectiveness (Yu et al., 2021) and is more efficient as it avoids the extra latency caused by the rewriting step.
@@ 2.1 Conversational Search
Currently, conversational search primarily relies on two main methods: conversational query rewriting (CQR) and conversational dense retrieval (CDR). CQR (Yu et al., 2020; Wu et al., 2022; Kumar and Callan, 2020; Voskarides et al., 2020; Lin et al., 2020; Mao et al., 2023b; Liu et al., 2021; Vakulenko et al., 2021a, b; Mao et al., 2023c; Mo et al., 2023a) transforms the whole session into a context-independent query. The generated query rewrites can directly perform ad-hoc retrieval. In contrast, CDR (Yu et al., 2021; Mo et al., 2024; Krasakis et al., 2022; Mao et al., 2022a; Mo et al., 2023b; Mao et al., 2023d, 2022b; Dai et al., 2022) aims to train a session encoder that is capable of encoding the conversational context into a high-dimensional space for conducting dense retrieval. However, the session embedding encoded by the conversational query encoder lacks interpretability, hindering developers from comprehending the retrieval results obtained by the search engine.
@@ 4.1 Datasets
We use four public conversational search datasets: QReCC (Anantha et al., 2021), TREC CAsT-19 (Dalton et al., 2020), TREC CAsT-20 (Dalton et al., 2021a), and TREC CAsT-21 (Dalton et al., 2021b). The QReCC dataset consists of 13.6K conversations, with an average of 6 turns per conversation. While the three CAsT datasets (19, 20, 21) only comprise 50, 25, and 26 conversations, respectively, but with more detailed relevance labeling. All four datasets provide human rewrites for each turn. Following existing works (Mao et al., 2023d; Mo et al., 2023a), we train CDR models on the QReCC dataset and conduct evaluations on the three CAsT datasets.
@@ 4.4 Evaluation Metrics
Retrieval and inversion evaluation. Following existing works (Mo et al., 2023a; Mao et al., 2023d) and the official settings of the CAsT datasets (Dalton et al., 2021a), we choose MRR, NDCG@3, and Recall@100 to evaluate the retrieval performance. We use two metrics to quantify the fidelity of the embedding inversion: (1) The absolute difference in the retrieval performances between using the session embeddings and the transformed text. (2) Following Vec2Text (Morris et al., 2023), we also calculate the cosine similarity between the session embeddings and the transformed text embeddings.
@@ 5.1 Retrieval and Inversion Evaluation
Method
CAsT-19 CAsT-20 CAsT-21
UniCRR
94.10 92.80 87.90
ConvInv
95.80 95.20 94.50

### v2

@@ Abstract
Conversational dense retrieval has shown to be effective in conversational search. However, a major limitation of conversational dense retrieval is their lack of interpretability, hindering intuitive understanding of model behaviors for targeted improvements. This paper presents ConvInv , a simple yet effective approach to shed light on interpretable conversational dense retrieval models. ConvInv transforms opaque conversational session embeddings into explicitly interpretable text while faithfully maintaining their original retrieval performance as much as possible. Such transformation is achieved by training a recently proposed Vec2Text model Morris et al. (2023) based on the ad-hoc query encoder, leveraging the fact that the session and query embeddings share the same space in existing conversational dense retrieval. To further enhance interpretability, we propose to incorporate external interpretable query rewrites into the transformation process. Extensive evaluations on three conversational search benchmarks demonstrate that ConvInv can yield more interpretable text and faithfully preserve original retrieval performance than baselines. Our work connects opaque session embeddings with transparent query rewriting, paving the way toward trustworthy conversational search. Our code is available at this repository.
@@
Yiruo Cheng, Kelong Mao, Zhicheng Douthanks: Corresponding author. Gaoling School of Artificial Intelligence, Renmin University of China {chengyr,mkl,dou}@ruc.edu.cn
@@ 1 Introduction
With the rapid development of language modeling, conversational search has emerged as a novel search paradigm and is garnering more and more attention. Different from the traditional ad-hoc search paradigm characterized by keyword-based queries and “ten-blue” links (Yu et al., 2020), conversational search empowers users to interact with the search engine through multi-turn natural language conversations to seek information, which brings a more intuitive and efficient search experience (Mao et al., 2022b; Gao et al., 2022; Zhu et al., 2023).
In conversational search, the system input is a multi-turn natural language conversation, which may have many linguistic problems such as omissions, co-references, and ambiguities (Radlinski and Craswell, 2017), posing great challenges for accurately grasping the user’s real information needs. Recently, conversational dense retrieval (CDR) (Yu et al., 2021; Lin et al., 2021; Kim and Kim, 2022; Mao et al., 2022a; Qian and Dou, 2022; Mo et al., 2023b; Chen et al., 2024), which directly encodes the whole conversational search session and the passages into a unified embedding space to perform matching, has shown to be a promising method to solve this complex search task. Compared to another type of method: conversational query rewriting (CQR) (Lin et al., 2020; Vakulenko et al., 2021a; Wu et al., 2022; Mo et al., 2023a), which is a two-step method that first reformulates the search session into a decontextualized query rewrite and subsequently inputs this rewrite into existing ad-hoc search models for search, the end-to-end CDR models can be directly optimized towards better search effectiveness (Yu et al., 2021) and is more efficient as it avoids the extra latency caused by the rewriting step.
@@ 2.1 Conversational Search
Currently, conversational search primarily relies on two main methods: conversational query rewriting (CQR) and conversational dense retrieval (CDR). CQR (Yu et al., 2020; Wu et al., 2022; Kumar and Callan, 2020; Voskarides et al., 2020; Lin et al., 2020; Mao et al., 2023b; Liu et al., 2021; Vakulenko et al., 2021a, b; Mao et al., 2023c; Mo et al., 2023a) transforms the whole session into a context-independent query. The generated query rewrites can directly perform ad-hoc retrieval. In contrast, CDR (Yu et al., 2021; Mo et al., 2024; Krasakis et al., 2022; Mao et al., 2022a; Mo et al., 2023b; Mao et al., 2023d, 2022b; Dai et al., 2022; Hai et al., 2023; Mao et al., 2024) aims to train a session encoder that is capable of encoding the conversational context into a high-dimensional space for conducting dense retrieval. However, the session embedding encoded by the conversational query encoder lacks interpretability, hindering developers from comprehending the retrieval results.
@@ 4.1 Datasets
We use four public conversational search datasets: QReCC (Anantha et al., 2021), TREC CAsT-19 (Dalton et al., 2020b), TREC CAsT-20 (Dalton et al., 2020a), and TREC CAsT-21 (Dalton et al., 2021). The QReCC dataset consists of 13.6K conversations, with an average of 6 turns per conversation. While the three CAsT datasets (19, 20, 21) only comprise 50, 25, and 26 conversations, respectively, but with more detailed relevance labeling. All four datasets provide human rewrites for each turn. Following existing works (Mao et al., 2023d; Mo et al., 2023a), we train CDR models on the QReCC dataset and conduct evaluations on the three CAsT datasets.
@@ 4.4 Evaluation Metrics
Retrieval and inversion evaluation. Following existing works (Mo et al., 2023a; Mao et al., 2023d) and the official settings of the CAsT datasets (Dalton et al., 2020a), we choose MRR, NDCG@3, and Recall@100 to evaluate the retrieval performance. We use two metrics to quantify the fidelity of the embedding inversion: (1) The absolute difference in the retrieval performances between using the session embeddings and the transformed text. (2) Following Vec2Text (Morris et al., 2023), we also calculate the cosine similarity between the session embeddings and the transformed text embeddings.
@@ 5.1 Retrieval and Inversion Evaluation
Method CAsT-19 CAsT-20 CAsT-21
UniCRR 94.10 92.80 87.90
ConvInv 95.80 95.20 94.50
@@ 6 Conclusion
Acknowledgments
This work was supported by National Key R&D Program of China No. 2022ZD0120103, National Natural Science Foundation of China No. 62272467, the fund for building world-class universities (disciplines) of Renmin University of China, and Public Computing Cloud, Renmin University of China. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.